Using the midpoint method, the elasticity of Cho's labor supply between the wages of $50 and $65 per hour is approximately

Calculating the price elasticity of supply 

Cho is a retired teacher who lives in Dallas and does some consulting work for extra cash. At a wage of SSO per hour, she is willing to work iO hours per week. At $65 per hour, she is willing to work 19 hours per week. 

Using the midpoint method, the elasticity of Cho's labor supply between the wages of $50 and $65 per hour is approximately _______ . which means that Cho's supply of labor over this wage range is _______ .
Answer #1

Elasticity of labour supply = Change (%) in labour hours supplied/ Change (%) in wage rate

Elasticity of labour supply = (L2 - L1)/(W2 - W1) x (W2 + W1)/(L2 + L1) = 9/15 x 115/29

Elasticity = 2.38

Cho's supply of labour over this wage range is elastic. (absolute value > 1)